# MedS3: Towards Medical Small Language Models with Self-Evolved Slow Thinking

### v1

@@ Abstract
Medical language models (MLMs) have become pivotal in advancing medical natural language processing. However, prior models that rely on pre-training or supervised fine-tuning often exhibit low data efficiency and limited practicality in real-world clinical applications. While OpenAI’s O1 highlights test-time scaling in mathematics, attempts to replicate this approach in medicine typically distill responses from GPT-series models to open-source models, focusing primarily on multiple-choice tasks. This strategy, though straightforward, neglects critical concerns like data privacy and realistic deployment in clinical settings. In this work, we present a deployable, small-scale medical language model, MedS3, designed for long-chain reasoning in clinical tasks using a self-evolution paradigm. Starting with a seed dataset of around 8,000 instances spanning five domains and 16 datasets, we prompt a base policy model to perform Monte Carlo Tree Search (MCTS) to construct verifiable reasoning chains. Each reasoning step is assigned an evolution rollout value, allowing verified trajectories to train the policy model and the reward model. During inference, the policy model generates multiple responses, and the reward model selects the one with the highest reward score. Experiments on eleven evaluation datasets demonstrate that MedS3 outperforms prior open-source models by 2 points, with the addition of the reward model further boosting performance ($\sim$13 points), surpassing GPT-4o-mini. Code and data are available at https://github.com/pixas/MedSSS.
@@
Shuyang Jiangthanks: Equal contribution.,♣,♢, Yusheng Liao∗,♠,♢, Zhe Chen♠,♢, Ya Zhang♠,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory shuyangjiang23@m.fudan.edu.cn {liao20160907,chenzhe2018,wangyanfeng622,ya_zhang,yuwangsjtu}@sjtu.edu.cn
@@ 1 Introduction
Models Without Without Step-wise Small Size Slow Thinking Process
Pretraining Close-sourced Teacher Search Support Reward Usage
@@
Large Language Models (LLMs) have demonstrated significant potential in the medical domain (Singhal et al., 2023; Nori et al., 2023; Chen et al., 2023b), with applications ranging from generating clinical notes(Biswas and Talukdar, 2024; Jung et al., 2024) to supporting patient communication (Tu et al., 2024; Liao et al., 2024b). Recently, slow-thinking reasoning models, exemplified by OpenAI o1, have shown impressive improvements on reasoning-intensive mathematical problems(OpenAI, 2024; Wang et al., 2024). However, limited efforts have been made to induce similarly strong reasoning abilities in medical-oriented language models.
@@
Before the rise of OpenAI o1, researchers focused on improving medical language models (MLMs) primarily through extensive pre-training, which demands large computational resources and often yields only modest gains in downstream tasks. For instance, MMed-Llama3 post-pretrained Llama3 8B on 25.5B tokens but achieved only about a five-point performance improvement compared to its base model (Qiu et al., 2024). Other works (Christophe et al., 2024; Ankit Pal, 2024) using supervised fine-tuning(Ouyang et al., 2022) are more computationally efficient but impart limited knowledge to the model, restricting performance gains on tasks outside the fine-tuning corpus. Moreover, many human-annotated medical datasets provide only concise responses (either short phrases or a single ground truth option), and fine-tuning on such data can degrade an MLM’s broader language fluency, reducing its practical value in realistic clinical scenarios. On the other hand, synthetic corpora generated by LLMs (Luo et al., 2024; Qiu et al., 2024) often contain hallucinations (Xu et al., 2024b; Huang et al., 2023), so directly using these outputs for teacher-forcing can limit a target model’s optimization space. Consequently, focusing on inference-time scaling, often referred to as “slow thinking,” emerges as a data-efficient approach with the potential to enhance general performance while mitigating the drawbacks of pre-training- or fine-tuning-heavy strategies.
@@
A closely related work to slow thinking is HuatuoGPT-o1 (Chen et al., 2024), which employs GPT-4o to generate reasoning-intensive problems and corresponding complex reasoning steps for distillation and reinforcement learning. While it achieves certain levels of long-chain reasoning, its heavy reliance on large proprietary models and the inherent instability of Proximal Policy Optimization (PPO) training (Dulac-Arnold et al., 2021) limits its generalizability to other clinical applications.
@@
Another relevant work, O1 Journey Part 3 (Huang et al., 2025), directly distills OpenAI o1’s outputs into $\sim$70B-parameter models without leveraging a fine-grained process reward model (PRM). This approach compromises user-friendliness and data privacy, and it also struggles to detect hallucinations. Moreover, relying exclusively on multiple-choice problems for distillation constrains its applicability to a broader range of clinical tasks.
@@
In contrast, this work introduces MedS3, a small-scale medical language model equipped with robust long-chain reasoning capabilities (the “policy”) and an accompanying fine-grained process reward model (the “PRM”). Our approach uses a self-bootstrapping pipeline to iteratively enhance the model’s performance across diverse clinical tasks. Specifically, we curate 16 medical tasks from established medical training corpora, encompassing clinical diagnosis QA, medical natural language inference, knowledge-intensive QA, long-context QA, and biomedical QA. We then sample 500 seed instances from each task to form an initial dataset of approximately 8,000 instances, which kickstarts our self-evolution process. Leveraging a Monte-Carlo Tree Search (MCTS)-based evolution pipeline, we iteratively generate synthetic datasets for both policy fine-tuning and PRM training. Verified synthetic trajectories are preserved to refine the policy model, while intermediate reasoning steps are labeled with MCTS rollout values to build the PRM dataset. By fine-tuning the base model on this enriched policy dataset and further enhancing it using the PRM, our method achieves state-of-the-art performance across eleven clinical benchmarks, surpassing GPT-4o-mini in tasks that demand precise verification of each reasoning step.
@@
The uniqueness of MedS3 is further underscored by a comprehensive comparison in Table 1, which demonstrates its superiority over other medical models in robust long-chain reasoning and breadth of clinical-task coverage.
Overall, we summarize our contributions as follows:
@@
    First Self-Evolution Framework: To the best of our knowledge, this work introduces the first self-evolution framework specifically designed to empower small-scale medical models with long-chain reasoning capabilities, enabling efficient performance across a wide range of clinical applications.
@@
    State-of-the-Art Performance: Our self-evolved system, MedS3, achieves state-of-the-art performance across eleven clinical benchmarks, outperforming all open-source models and GPT-4o-mini. This is driven by the integration of a fine-grained PRM that enhances reasoning accuracy at each step.
@@
    Open-Source Resources for Research: We openly release both the policy fine-tuning corpus and the process reward model corpus, providing valuable resources for future research and fostering further advancements in medical AI.
@@ 2 MedS3
Framework This section presents a detailed overview of the proposed MedS3 framework, which is structured into four key components:
@@
    Self-Bootstrapping Evolution which synthesizes reasoning trajectories for instances of the training set with Monte-Carlo Tree Search (MCTS) technique using the base policy $\pi_{0}$.
@@
    Policy Model $\pi$ which is derived by fine-tuning the base policy $\pi_{0}$ using the generated synthetic data in a supervised learning manner.
@@
    Process Reward Model (PRM) $V_{\theta}$ which is fine-tuned with step-wise supervision and assigns a float value in the range $[0,1]$ to each reasoning step, representing the probability of step-wise correctness.
@@
    PRM-guided Inference which utilizes the MCTS algorithm to generate diverse solutions and determines the final solution by evaluating them with the PRM.
@@
Figure 1: Overview of MedS3 framework. MedS3 utilizes a Monte-Carlo Tree Search pipeline to self-generate step-by-step reasoning paths for each question in the seed dataset (a). During this process, MedS3 uses result simulation to obtain the rollout value for each node (b); After obtaining the child’s rollout value, MedS3 executes back-propagate to enable precise value prediction from deeper layer to transfer back to shallow nodes (c). After gathering all correct and wrong finish nodes, we use supervised fine-tuning to optimize the policy model $\pi$ with correct reasoning trajectories and step-wise discriminative loss to obtain a process reward model $V_{\theta}$ (d).
@@ 2.1 MCTS-guided Evolution
This algorithm builds upon an $n$-ary tree, where each tree node $T$ contains the following attributes: (1) reasoning step $s$, which is an intermediate step tracing from the root; (2) value $v$, which is an evaluation of current contribution to the correctness of the whole reasoning trace; (3) the children nodes $\{c\}$, which is a collection of nodes that continue reasoning from the current node; (4) the parent node $p$ which is the former reasoning step and (5) the number of visits $n$. Every root node is initialized as $T=([s_{0}],0,\emptyset,\mathrm{null},0)$ where $s_{0}=$“Let s break down this problem step by step.” to guarantee a multi-step reasoning process. There are five stages in a full MCTS pipeline, including Node Selection, Node Expansion, Node Rollout, and Backpropagation.
@@
$UCB(T)=v_{C}+c\sqrt{\frac{\ln n_{T_{parent}}}{n_{T}}}$ (1)
where $T_{parent}$ is the parent of current node $T$ and $c$ is an exploration constant, which is set to 2. For each intermediate node, we select its child node with the highest $UCB$ value. We choose this criterion to expect models to further explore those nodes with high rollout values but low visiting counts, which inhibits models from repeatedly expanding already high-value nodes and encourages the expansion of underexplored nodes with the second highest value.
@@
After picking up the candidate node $T_{c}$ using the UCB criterion, we expand the reasoning steps of the current node. If the current node possesses a relatively high value ($v_{c}\geq thr$, where $thr=0.9$ is a pre-defined threshold), we prompt the node to directly generate a Finish node to accomplish this path reasoning. This manual operation not only does not impact the reasoning correctness as a value is close to 1 only when the trace $[s_{0},s_{1},\cdots,s_{k}]$ is close to the correct final answer, but also reduces unnecessary exploration tokens. Otherwise, assume that the selected node is located at $k$-th depth among the tree with previous reasoning trajectories $[s_{0},s_{1},\cdots,s_{k}]$ connected by a coherence phrase $t_{s}$, we sample $B$ single-step outputs $\{s_{k+1,i}\mid i=1,2,\cdots,B\}$ based on the previous trajectory using a Reason [^1] Prompts of Finish and Reason actions are illustrated in Appendix A node:
@@
As the PRM is not yet available during bootstrapping, the rollout process is mainly conducted using simulation to obtain the estimated value for a chosen node. The process is important since an accurate value estimation is not only crucial for the PRM training but also impacts the priority of node expansion in future exploration. To this end, for a chosen unvisited node $T_{c}$, we set a simulation budget $L=\min(L_{\min},\frac{L_{0}}{k})$ where $k$ is reasoning step counts of $T_{c}$, to encourage sufficient simulation trials when the known reasoning path is short, but expect to see an accurate reasoning result when the trajectory is long enough. After determining the simulation attempt counts, we prompt the policy model $\pi_{0}$ to directly output the answer $L$ times conditioning on the reasoning trace:
@@
where $l\in[1,L]$ and $a_{c}^{l}$ is the $l$-th simulated answer. The accuracy of the $L$ simulations $acc=\frac{1}{L}\sum_{l=1}^{L}\mathds{1}_{a_{c}^{l}=y}$ is assigned as the value of $T_{c}$. For the Finish node, the rollout value is computed by comparing the derived answer and the ground truth and an $0/1$ value is assigned. In terms of inference, the value is obtained by calling the PRM. We elucidate this process in §2.4.
@@
After obtaining the value for the selected node, we conduct value back-propagation starting from $T_{c}$ till the root node, updating all tree node values among the trace. This process aims to use more accurate value estimation in the deeper layer to update early nodes’ values so that these values serving as PRM’s optimization labels become synchronously precise. Specifically, for the node among the trace from the root node to $T_{k}$, we update its visits $n_{k}$ and $v_{k}$ as follows:
@@
Note that this update process considers not only the child value, but also incorporates the value of itself. This operation guarantees that the parent node will never have the same value as its children, which considers both correctness and completeness for the evaluation of a reasoning step.
@@
Due to computational limits, we cannot afford excessive computing resources to fully expand the tree for a training instance. Therefore, for a single reasoning trajectory, apart from the circumstance that the node is explicitly assigned as Finish due to a relatively high value, we propose to prompt the policy model $\pi_{0}$ to self-generate the Finish signal. Specifically, we prompt the model to determine which actions to take as the next step in the Node Expansion stage. For the $B$ children, we first prompt the model to generate $B$ actions from [Reason,Finish]. If the model is deterministic about the reasoning process, it opts to generate consistent Finish actions. Generation of more than one Reason actions indicates that the policy model is still of low confidence about the termination of reasoning. Otherwise, we also assign a Finish action for the expanded node and take the corresponding prompt template for thought generation. A finish node will never participate in the later node expansion so the search costs are reduced. The whole tree is terminated once the number of Finish nodes reach a pre-defined node limit.

### v2

@@ Abstract
Medical language models (MLMs) have become pivotal in advancing medical natural language processing. However, prior models that rely on pre-training or supervised fine-tuning often exhibit low data efficiency and limited practicality in real-world clinical applications. While OpenAI’s o1 highlights test-time scaling in mathematics, attempts to replicate this approach in medicine typically distill responses from GPT-series models to open-source models, focusing primarily on multiple-choice tasks. This strategy, though straightforward, neglects critical concerns like data privacy and realistic deployment in clinical settings. In this work, we present a deployable, small-scale medical reasoning system, MedS3, designed for long-chain reasoning in clinical tasks using a self-evolution paradigm. Starting with a seed dataset of around 8,000 instances spanning five domains and 16 datasets, we prompt a base policy model to perform Monte Carlo Tree Search (MCTS) to construct rule-verifiable reasoning chains. Each reasoning step is assigned an evolution rollout value, allowing verified trajectories to train the policy model and the process reward model (PRM). During inference, the policy model generates multiple responses, and the reward model selects the one with a newly proposed PRM-guided Vote-Sum (P-VS) strategy. Experiments on eleven evaluation datasets demonstrate that MedS3 outperforms not only the prior strongest medical model by 6.59, but also 32B-level general reasoning models by 8.71 points. Code and data are available at https://github.com/pixas/MedSSS.
@@
Shuyang Jiang♣,♢, Yusheng Liao♠,♢, Zhe Chen♠,♢, Ya Zhang♠,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory shuyangjiang23@m.fudan.edu.cn {liao20160907,chenzhe2018,ya_zhang,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn
@@ 1 Introduction
Models Without Without Diverse Small Size Slow Thinking Process
Pretraining Close-sourced Teacher Clinical Coverage Reward Usage
@@
Large Language Models (LLMs) have demonstrated significant potential in the medical domain (Singhal et al., 2023; Nori et al., 2023; Chen et al., 2023b), with applications ranging from generating clinical notes (Biswas and Talukdar, 2024; Jung et al., 2024) to supporting patient communication (Tu et al., 2024; Liao et al., 2024b). Recently, slow-thinking reasoning models, exemplified by OpenAI o1 (OpenAI, 2024), have shown impressive improvements on reasoning-intensive mathematical problems (Lyu et al., 2025; Wang et al., 2024). However, limited efforts have been made to induce similarly strong reasoning abilities in medical-oriented language models.
@@
Before the rise of OpenAI o1, researchers focused on improving medical language models (MLMs) primarily through extensive pre-training, which demands large computational resources and often yields only modest gains in downstream tasks. For instance, MMed-Llama3 post-pretrained Llama3 8B on 25.5B tokens but achieved only about a five-point performance gain compared to its base model (Qiu et al., 2024). Other works (Christophe et al., 2024; Ankit Pal, 2024) leverage supervised fine-tuning (SFT; Ouyang et al. (2022)) since it is computationally efficient. However, many human-annotated SFT datasets provide only concise responses (either short phrases or a single ground truth option), and hence fine-tuning on such data can degrade an MLM’s broader language fluency, reducing its practical value in realistic clinical scenarios. On the other hand, synthetic corpora generated by LLMs (Luo et al., 2024; Qiu et al., 2024) often contain hallucinations (Xu et al., 2024b; Huang et al., 2023), so directly using these outputs for teacher-forcing limits a target model’s optimization space. Consequently, focusing on inference-time scaling, often referred to as “slow thinking”, emerges as a data-efficient approach with the potential to enhance clinical mastery while mitigating the drawbacks of pre-training- or fine-tuning-heavy strategies.
@@
A closely related work to slow thinking is HuatuoGPT-o1 (Chen et al., 2024), which employs GPT-4o to generate reasoning-intensive problems and corresponding complex reasoning steps for distillation and reinforcement learning (RL). While it achieves certain levels of long-chain reasoning, its heavy reliance on large proprietary models limits its generalizability to other clinical applications. Moreover, The adopted RL algorithm (Schulman et al., 2017) requires accurate value models, which is far from practicability in medicine yet.
@@
Another relevant work, O1 Journey Part 3 (Huang et al., 2025), directly distills OpenAI o1’s outputs into $\sim$70B-parameter models. This approach compromises user-friendliness and data privacy, and also struggles to detect hallucinations. Moreover, relying exclusively on multiple-choice problems for distillation constrains its applicability to a broader range of clinical tasks.
@@
In contrast, this work introduces MedS3, a small-scale medical language model equipped with robust long-chain reasoning capabilities (the “policy”) and an accompanying fine-grained process reward model (the “PRM”). Our approach uses a self-bootstrapping pipeline to enhance the model’s performance across diverse clinical tasks. Specifically, we curate 16 medical tasks from established medical training corpora, encompassing clinical diagnosis QA, medical natural language inference, knowledge-intensive QA, long-context QA, and biomedical QA. We then sample 500 seed instances from each task to form an initial dataset of approximately 8,000 instances, which kickstarts our self-evolution process. Leveraging a Monte-Carlo Tree Search (MCTS)-based evolution pipeline, we simultaneously generate synthetic datasets for both policy fine-tuning and PRM training. Rule-verified synthetic trajectories are preserved to optimize the policy model, where intermediate reasoning steps labeled with MCTS rollout values serve as the PRM dataset. By fine-tuning the base model on this enriched policy dataset and further enhancing it using our proposed PRM-guided Vote-Sum strategy, our system achieves holistically superior performance across eleven clinical reasoning benchmarks, surpassing both medical models and general reasoning models with much lower parameters.
@@
The uniqueness of MedS3 is further underscored by a comprehensive comparison in Table 1, which demonstrates its superiority over other medical models in robust long-chain reasoning and breadth of clinical-task coverage. Overall, we summarize our contributions as follows:
@@
    First Self-Evolution Framework: To the best of our knowledge, this work introduces the first self-evolution framework specifically designed to empower small-scale medical models with long-chain reasoning capabilities, enabling data-efficient performance gains across a wide range of clinical applications.
@@
    State-of-the-Art Performance: Our self-evolved system, MedS3, achieves comprehensive performance improvements across eleven clinical benchmarks, outperforming all medical and large-sized reasoning models overall. This is driven by the integration of a fine-grained PRM that enhances reasoning accuracy at each step.
@@
    Open-Source Resources for Research: We openly release both the policy fine-tuning corpus and process reward model corpus, providing valuable resources for future research and fostering further advancements in medical AI.
@@ 2 MedS3
This section presents a detailed overview of the proposed MedS3 framework, which is structured into four key components:
@@
    Self-Bootstrapping Evolution (§2.1) which synthesizes reasoning trajectories as training data, with Monte-Carlo Tree Search (MCTS) technique using the base policy $\pi_{0}$.
@@
    Policy Model $\pi$ (§2.2) which is derived by fine-tuning the base policy $\pi_{0}$ using the generated synthetic data with supervised learning.
@@
    Process Reward Model (PRM) $V_{\theta}$ (§2.3) which is fine-tuned with step-wise supervision using soft-labels and assigns a value in the range $[0,1]$ to each reasoning step, representing correctness of individual steps.
@@
    PRM-guided Inference (§2.4) which utilizes the proposed PRM-guided Vote-Sum strategy to choose the final solution.
@@
Figure 1: Overview of the construction of MedS3 framework. MedS3 utilizes a Monte-Carlo Tree Search pipeline to self-generate step-by-step reasoning paths for each question in the seed dataset (a). During this process, MedS3 uses result simulation to obtain the rollout value for each node (b); After obtaining the child’s rollout value, MedS3 executes back-propagate to enable precise value prediction from deeper layer to transfer back to shallow nodes (c). After gathering all correct and wrong finish nodes, we use supervised fine-tuning to optimize the policy model $\pi$ with correct reasoning trajectories and step-wise discriminative loss to obtain a process reward model $V_{\theta}$ (d).
@@ 2.1 MCTS-guided Evolution
This algorithm builds upon an $n$-ary tree, where each tree node $T$ contains the following attributes: (1) reasoning step $s$, which is an intermediate step tracing from the root; (2) value $v$, which is an evaluation of the correctness of the current node; (3) the children nodes $\{c\}$, which is a collection of nodes that continue reasoning from the current node; (4) the parent node $p$ which is the former reasoning step and (5) the number of visits $n$. Every root node is initialized as $T=([s_{0}],0,\emptyset,\mathrm{null},0)$ where $s_{0}=$“Let s break down this problem step by step.” to guarantee a multi-step reasoning process. There are four stages in a full MCTS pipeline, including Node Selection, Node Expansion, Node Rollout, and Backpropagation.
@@
$UCB(T)=v_{C}+\gamma\sqrt{\frac{\ln n_{T_{parent}}}{n_{T}}}$ (1)
where $T_{parent}$ is the parent node of $T$ and $\gamma$ is an exploration constant set as 2. For each intermediate node, we select its child node with the highest $UCB$ value. We choose this criterion to expect models to further explore those nodes with high rollout values but low visiting counts, which inhibits models from repeatedly expanding already high-value nodes and encourages the expansion of underexplored nodes with the second highest value.
@@
After picking up the candidate node $T_{c}$ using the UCB criterion, we expand the reasoning steps of the current node. If the current node possesses a relatively high value ($v_{c}\geq thr$, where $thr=0.9$ is a pre-defined threshold), we prompt the node to directly generate a Finish node to accomplish this path reasoning. This manual operation not only does not impact the reasoning correctness as a value is close to 1 only when the trace $[s_{0},s_{1},\cdots,s_{k}]$ is close to the correct final answer, but also reduces unnecessary exploration tokens. Otherwise, assume that the selected node is located at $k$-th depth among the tree with previous reasoning trajectories $[s_{0},s_{1},\cdots,s_{k}]$ connected by a coherence phrase $t_{s}$, we sample $B$ single-step outputs $\{s_{k+1,i}\mid i=1,2,\cdots,B\}$ based on the previous trajectory using a Reason [^1] Prompts of Finish and Reason actions are illustrated in Appendix D node:
@@
As the PRM is not yet available during evolution, the rollout process is conducted using simulation to obtain the estimated value for a chosen node. Specifically, for a chosen unvisited node $T_{c}$, we set a simulation budget $L=\min(L_{\min},\frac{L_{0}}{k})$ where $k$ is reasoning step counts, to encourage sufficient simulation trials when the known reasoning path is short, but expect to see a deterministic reasoning result conditioning on a long trajectory. After that, we prompt the policy model $\pi_{0}$ to directly output the answer $L$ times:
@@
where $l\in[1,L]$ and $a_{c}^{l}$ is the $l$-th simulated answer. The accuracy of the $L$ simulations $acc=\frac{1}{L}\sum_{l=1}^{L}\mathds{1}_{a_{c}^{l}=y}$ is assigned as the value of $T_{c}$. For the Finish node, the rollout value is computed by comparing the derived answer and the ground truth and an $0/1$ value is assigned.
@@
After obtaining the value for the selected node, we conduct value back-propagation starting from $T_{c}$ till the root node, updating all tree node values among the trace. This process aims to use more accurate value estimation in the deeper layer to update early nodes’ values so that these values serving as PRM’s optimization labels become synchronously precise. Specifically, for an arbitrary node $T_{k}$, we propose to update its visits $n_{k}$ and $v_{k}$ as follows:
@@
Note that this update process considers both its children values and its own value, which guarantees that the parent node will never have the same value as its children. It considers both correctness and completeness for the evaluation of a reasoning step.
@@
Due to computational limits, we cannot afford excessive computing resources to fully expand the tree for a training instance. Therefore, we set two criteria to terminate the search process. Firstly, we set the minimum correct nodes to $\tau$. Once the total correct count in the tree exceeds $\tau$, we stop the exploration of this tree. Second, if there are no correct nodes after affording a certain number of node exploration trials, we prompt $\pi_{0}$ to generate Finish node for all leaves. The first strategy aims to reduce extra computation, while the second expects to obtain as more correct answers as possible, for the optimization of both policy and reward models.